OpenStudy (anonymous):

L(0, -8), M(4, 3) Whats the midpoint.? could some one please help me.

OpenStudy (anonymous):

Use the midpoint formula! (x1 + x2)/2 (y1+y2)/2

OpenStudy (isaiah.feynman):

Midpoint formula. \[(\frac{ x_{1}+x_{2} }{ 2 },\frac{ y_{1}+y_{2} }{ 2 })\]

OpenStudy (anonymous):

After the what should i do next.?

OpenStudy (isaiah.feynman):

Substitute!!!

OpenStudy (anonymous):

okay then so i got 34.?

OpenStudy (anonymous):

\[\sqrt{34}\]

OpenStudy (isaiah.feynman):

I assume you are right cos that stuff is easy, you can't make mistakes.

OpenStudy (anonymous):

Okay thank you :D so im right .?

OpenStudy (isaiah.feynman):

Oh no your wrong its (2,-5/2)

OpenStudy (anonymous):

What how .?

OpenStudy (anonymous):

im sorry may you please help me out.?

OpenStudy (anonymous):

lol. Oh boy okay! (2, -5/2)

OpenStudy (anonymous):

ahaha we where both wrong .

OpenStudy (anonymous):

The answer is (2,-5/2) not (4,-5/2) because 4/2 doesnt equal 4 xD

OpenStudy (isaiah.feynman):

Lol whats wrong with me?

OpenStudy (anonymous):

Hey! we all make mistakes right? :)

OpenStudy (anonymous):

lol yea we do ahah :D and Isaiah because you first said the right answer and then changed it lol :P